### v1

Abstract
[…]
keywords:
Generalization, Source-free Domain Adaptation, Large Vision Foundation Models, Object Detection
1 Introduction
Medical object detection is a well-studied problem in computer vision [10, 32, 153, 141, 129]. The success of deep learning techniques for the problem has been supported by the abundance of extensively annotated detection datasets [23, 33, 39, 76, 134], which facilitates the supervised training of deep object detection models. Unsupervised Domain Adaptation (UDA) offers a strategy for adapting object detectors to new domains where labeled data is not available.
This arises from the growing reliance on medical data, where annotations are not only expensive to obtain but also prone to errors in complex or ambiguous scenarios. It has been widely observed that deep object detection models often struggle to generalize to new visual domains despite their effectiveness in familiar visual contexts. UDA is a popular solution strategy [16, 37, 47, 48, 81, 100, 107], which bridges the gap between source and target domains by aligning the feature distributions [17, 46, 54, 98, 110, 53, 147]. However, the strategy requires access to source-domain data at the adaptation stage, which severely limits its applicability in medical imaging [74, 80, 122]. This motivates us to focus on Source-Free Object Detection (SFOD) in this work.
SFOD has received significant attention for the image classification task in recent years [130, 143, 120, 57, 31, 51, 131]. However, there are relatively fewer works specifically addressing SFOD [51, 111, 88, 77, 21]. Given the complexities of cluttered backgrounds, viewpoint variations, and large number of negative samples in medical images, directly applying conventional SFOD methods for classification tasks to object detection in medical imaging leads to unsatisfactory performance. As these methods ignore the critical issue of context imbalance, which is a problem in many real-life scenarios, such as medical imaging. Thus, there is a need to develop SFOD techniques specially tailored for medical imaging due to clinical relevance [6].
Moreover many popular SFOD techniques utilize a self-supervised approach in a student-teacher (ST) framework. These methods bootstrap by training on pseudo-labels generated by a source-pretrained model  [67, 123, 70, 21, 112]. However, if the source data is biased, or the domain shift between source and target domains is significant, there is noise in the pseudo-labels, which impacts the training of a student model [27]. Since the pseudo-label error is significant, the, Exponential Moving Average (EMA) step, which updates the teacher model from the student model’s weight, ends up corrupting the teacher model as well leading to a mode collapse [3]. This phenomenon of progressive degradation, especially in source-free setting where source data is not available during adaptation, is visually evident in Figure 1, where traditional Mean Teacher frameworks and SF-Student-Teacher frameworks suffer from declining detection quality over time, while our method remains robust over the epochs. This is typically not a problem in the UDA setting, as supervised data from the source domain, acts as an anchor and prevents error accumulation in the student. Then the EMA step ensures that the teacher model does not get corrupted at any point in the adaptation.
Figure 1: Effect of pseudo-label noise and EMA updates on predictions across epochs for different SFOD methods (ST, SF-ST, and Ours). ST refers to Mean Teacher [105] and SF-ST refers to source-free setting. Both exhibit progressive degradation, while our proposed GT method preserves accurate and consistent performance over time.
To address these challenges, in this paper, we introduce our Grounded Teacher (GT), specifically designed to tackle context imbalance and mode collapse in the SFOD setting. Our approach introduces a Relation-Contextual Module (RCM) that explicitly models class biases and inter-class dynamics within the training data. Leveraging these insights, we implement a semantic augmentation strategy that enhances minority class representation by strategically blending them with visually similar majority class instances, drawn from our Cropbank repository[140]. This augmentation operates bidirectionally across domains, simultaneously addressing both class imbalance and domain shift. To further mitigate inter-class bias, we propose a Semantic-Aware Loss that dynamically prioritizes minority classes during training, particularly focusing on challenging cases where they are most vulnerable to misclassification as majority classes. Together, these components form a comprehensive framework that holistically addresses representation learning challenges in cross-domain scenarios.
Another key insight of this paper is that one can use learnt representations from Vision Foundational Models to guide a student model in SFOD settings. Due to the excellent zero shot capabilities of the Vision Foundation Models (VFMs), they can effectively guide student models to learn the right representation, and distances between the samples, even when the target domain has a large shift from the source domain. This helps the model overcome mode collapse caused by biased pseudo-labels during unsupervised training.
Through the integration of these approaches, we observe enhanced pseudo-label quality, leading to measurable gains on medical imaging benchmarks achieving a new state-of-the-art (SOTA) of 50.8 mAP in natural images for Cityscapes $\to$ Foggy Cityscapes, beating the previous standard of 45.0 mAP by +5.8 mAP ($\sim$13performance gain). We summarize the contributions of this paper as follows:
  • We highlight the problem of training instability of a student model in SFOD setting due to incorrect pseudo labels from a source pretrained model in the presence of source bias, or large domain shift between source and target domains.
  • We propose our GT model, supported by our RCM, which can map the model’s existing class biases. To this end, we formulate a Semantic Aware Loss (SAL) to further enhance the performance of biased classes.
  • We propose to leverage the zero-shot performance of VFMs to guide the student model in a sample-specific manner and help learn the correct representation in a large domain shift scenario.
  • We observe that SFOD formulation is especially used for medical imaging scenarios where sharing source data is harder due to stricter privacy constraints, and where source data bias is more likely due to smaller datasets captured in a single center mode. Hence, we experiment and demonstrate SOTA results on cross-domain datasets for Breast Cancer Detection. We report a recall of 0.45 and 0.43 at an FPI of 0.3 on publicly available DDSM to INBreast [85] and DDSM  to RSNA datasets respectively against the performance of 0.05 and 0.31 by the current SOTA [110].
2 Related Work
Object Detection. In recent years, object detection has become a central task in computer vision, garnering significant attention due to its wide range of applications [19, 154]. Traditional object detectors typically perform bounding box regression and category classification using different paradigms such as anchors [92, 75, 93], region proposals [94, 40], and point-based methods [115, 151, 106]. These approaches often rely on large-scale datasets like COCO [76] and PASCAL [33], which require extensive manual annotation and substantial data per object category. To reduce annotation costs, Weakly Supervised Object Detection (WSOD) methods have emerged as a promising alternative [139, 138]. These methods utilize image-level labels [83] instead of instance-level annotations, treating each image as a bag of proposals and leveraging Multiple Instance Learning (MIL) [7, 22, 8, 119] to assign labels to regions of interest. This cost-effective approach enables expansion of the detection vocabulary using only classification-level data [26, 150, 119, 65, 42, 113]. More recently, transformer-based models [4] have gained prominence in object detection due to their ability to model token-wise dependencies and contextual relationships. Inspired by the Vision Transformer (ViT) [32] and Detection Transformer (DETR) [10], several transformer-based detectors have been proposed [10, 153, 149, 25, 141, 129, 5], offering state-of-the-art performance across benchmarks. In biomedical imaging, detection techniques are increasingly adopted for critical tasks such as cancer diagnosis. Varlamova et al. [108] proposed MAMM-Net, based on Mask2Former, for fusing dual-view mammography features to localize malignant lesions, mimicking a radiologist’s diagnostic reasoning. Wang et al. [118] utilized an anchor-free FCOS network with segmentation-guided enhancement for lesion detection in ultrasound images. Alashban [2] employed a CBAM-enhanced YOLOv5 for breast cancer detection in Digital Breast Tomosynthesis (DBT) images, while Kebede et al. [59] also used YOLOv5 to identify suspicious masses in mammograms, emphasizing interpretability and diagnostic assistance. Yan et al. [128] extended Mask R-CNN to detect lesions in CT scans using bounding boxes. Despite these advances, most approaches remain reliant on supervised learning. In this work, we aim to improve model adaptability across domains without requiring additional annotations.
Unsupervised Domain Adaptation (UDA). UDA seeks to reduce the gap between domains by utilizing unlabeled target domain samples in object detection tasks. It is particularly effective when labeled data is accessible in the source domain for adaptation. Chen et al. [17] pioneered DAOD, exploring adversarial feature alignment for Faster R-CNN (FRCNN) [95]. Subsequent works [46, 15, 63, 14, 13, 96, 125] have investigated various aspects of feature alignment, while others [114, 53, 41, 147] have elaborately designed alignments for transformer architectures. Among these methodologies, approaches employing the student-teacher (ST) framework have gained prominence due to their superior performance in experimental evaluations. For example, Li et al. [71] employed adversarial alignment and weak-strong augmentation techniques to reduce the false positive ratio of pseudo-labels, while He et al. [44] mitigated domain shift by incorporating target object knowledge through self-distillation. Kennerley et al. [60] introduced the Class-Aware Teacher (CAT) framework, employing an Inter-Class Relation module (ICRm) to model class biases and using Class-Relation Augmentation (CRA) with an Inter-Class Loss (ICL) to mitigate class imbalance, demonstrating improved performance on UDA benchmarks. Additionally, Zhao et al. [147] utilized mask image modeling to align source and target domains and learn domain-invariant features. UDA has also been explored for medical image applications. Ryan et al. [97] used adversarial UDA with a U-Net architecture to segment breast tissue in mammograms from different acquisition devices. UDA has also been used for cell detection in digital pathology, where Pina et al. [89] utilized adversarial feature alignment with detection transformers to address cross-stain generalization in breast cancer analysis. It was also actively explored for breast cancer detection from mammograms, as demonstrated by Ashraf et al. [6] who introduced D-MASTER, a transformer-based ST framework with adaptive masking and confidence refinement. UDA has also been extended to leverage multiple source domains, particularly in medical image analysis, as seen in the work by Yang et al. [132] who introduced BTMuda, a bi-level multi-source UDA framework for breast cancer diagnosis using mammograms. Despite their significant progress, these methods heavily rely on access to source data, which can be challenging to obtain in real-life scenarios due to privacy concerns and data constraints. In the absence of source data, these methods encounter challenges with low-quality pseudo-labels, often resulting in a limited number of bounding boxes and false positives. To address these issues, we build upon the baseline of the ST framework with Exponential Moving Average (EMA), introducing a foundation-based Expert branch and semantic-aware loss to address these issues.
Source-Free Object Detection (SFOD). In real-world applications, access to source data during the adaptation phase is often limited due to privacy concerns, data transmission limitations, or proprietary restrictions, especially in medical imaging. SFOD refers to the approach of adapting a pre-trained source model to a target domain without needing access to the original source data [82]. Due to these constraints, SFOD has become more challenging compared to traditional Unsupervised Domain Adaptation (UDA) methods, leading to the emergence of SFOD as a distinct branch within the domain adaptation framework in recent years. Given the intricate nature of object detection tasks, which involve numerous regions, multi-scale features, and complex network architectures, coupled with the absence of source data and target pseudo-labels, it is evident that the straightforward application of existing UDA methods is inadequate. The proposed SFOD in [70] uses self-entropy descent to create high-quality pseudo-labels for self-training. SOAP [123] introduces domain perturbation on the target data, aiding the model in acquiring domain-invariant features resilient to the perturbations. LODS [67] introduces a style enhancement module and a graph alignment constraint to facilitate the model in learning domain-independent features. AASFOD [21] categorizes target images into source-similar and source-dissimilar images, followed by employing adversarial alignment with ST models. IRG [110] introduces an instance relation graph network integrated with contrastive loss to facilitate contrastive representation learning. PETS [77] introduces a dynamic teacher within the ST framework and proposes a consensus mechanism to integrate predictions from both the static teacher and the dynamic teacher models. While works in medical imaging have not explicitly used SFOD, they have utilized techniques that reflect the fundamental ideals of SFOD like Yan et al. [128] use Mask R-CNN and pseudo-masks, Akkus et al. [1] discuss transfer learning in segmentation, Wang et al. [117] use auxiliary segmentation, various works [59, 2, 61] employ advanced detection architectures and transfer learning for medical image analysis. These approaches highlight the principles of adapting models for detection tasks in new contexts, mirroring the goals of SFOD. Despite leveraging the ST framework, most of these approaches often overlook the challenge of overcoming local optima and training instability resulting from a single updated teacher model. To address this, we propose a Grounded Teacher framework. It leverages knowledge from expert models to mitigate catastrophic forgetting and uses relation contextual module for generating refined pseudo-labels by removing context bias.
3 Preliminaries
Problem statement. Let $X$ denotes the input space and $\mathcal{D}_{s}=\{(x_{s}^{i},y_{s}^{i})\}_{i=1}^{N_{s}}$, the source domain, where $y_{s}^{i}=(\mathbf{b_{s}^{i}},c_{s}^{i})$ is a tuple comprising of ground truth bounding box ($\mathbf{b_{s}^{i}}\in\mathbf{\mathcal{R}^{4}}$) and the corresponding class label ($c_{s}^{i}\in\mathbf{\mathcal{R}}$) of the object respectively. Here, $x_{s}\in X$ denotes the image from the source domain, and $N_{s}$ denotes the total number of source images. The target domain $\mathcal{D}_{t}=\{x_{t}^{i}\}_{i=1}^{N_{t}}$ is unlabeled, where $N_{t}$ denotes the total number of target images, and the target samples $x_{t}^{i}\in X$. Our objective is to transfer the source-trained model to the target domain without accessing samples from the source domain.
Mean Teacher Framework. The self-supervised adaptation strategy involves updating the student model using unlabeled target data, leveraging pseudo-labels generated from the teacher model. Both the student and the teacher models are initialized by the source-trained model. The pseudo labels undergo a confidence threshold-based filtering process, and only the reliable ones are utilized to supervise the student training [62]. The pseudo-label supervision loss for the object detection model can be expressed as:
$\mathcal{L}_{\text{stu}}=\mathcal{L}_{\text{box}}^{S}(x_{t}^{i},yt_{t}^{i})+\mathcal{L}_{\text{giou}}^{S}(x_{t}^{i},yt_{t}^{i})+\mathcal{L}_{\text{cls}}^{S}(x_{t}^{i},yt_{t}^{i}).$ (1)
Here $yt_{t}^{i}$ represents the pseudo-label for image $i$ obtained after filtering low-confidence predictions from the teacher. The teacher is updated progressively through an Exponential Moving Average (EMA) of student weights. We implement a hard threshold, $\tau$, on the classification scores generated by the teacher to ensure that only pseudo-labels with high confidence are utilized by the student network, thereby encouraging more reliable learning outcomes.
Thus, the overall self-training process for the traditional student-teacher (ST) based object detection framework is expressed as follows:
$\Theta_{s}$ $\leftarrow\Theta_{s}+\gamma\frac{\partial L_{\text{stu}}}{\partial\Theta_{s}},$ (2)
$\Theta_{t}$ $\leftarrow\alpha\Theta_{t}+(1-\alpha)\Theta_{s}.$ (3)
Here $\mathcal{L}_{\text{stu}}$ represents the student loss computed using pseudo-labels from the teacher network, and $\Theta_{s}$ and $\Theta_{t}$ denote student and teacher network, respectively. The parameters $\gamma$ and $\alpha$ denote the student learning rate and teacher EMA rate, respectively. Although the ST framework enables knowledge distillation using noisy pseudo-labels, it alone is insufficient for learning high-quality target features in a source-free setting and leads to model degradation. Therefore, to enhance the features in the target domain, we introduce Visual Foundation Models (VFMs) as Expert and incorporate supervised loss for knowledge transfer from expert VFMs. Following [72], a discriminator is added to encourage domain invariant feature representations with an associated loss, $\mathcal{L}_{\rm dis}$.
Figure 2: Architecture of proposed Grounded Teacher (GT) Framework
4 Methodology
Our proposed method, Grounded Teacher (GT), illustrated in Figure 3, extends the standard mean-teacher framework by introducing a novel Relation Contextual Module (RCM). At the heart of GT, the RCM is designed to systematically capture and quantify class-specific biases in the model. Unlike conventional methods that address class imbalance in a general way, RCM explicitly models the semantic relationships between classes, with a particular focus on minority classes that are frequently misclassified as dominant ones.
This is achieved through the construction of a batch-wise confusion matrix based on pseudo-labels, which is normalized and continuously aggregated into a global matrix. This evolving matrix provides a robust and dynamic representation of the model’s bias landscape, forming the foundation for two key components of our method: Class-Relation Augmentation and Semantic Loss. Class-Relation Augmentation targets image-level class imbalance. Using the class relationships encoded by RCM, we identify minority classes in an image that share strong semantic overlap with dominant classes. To increase their representation and improve discrimination, we apply MixUp [142]—a technique known to be effective in imbalanced scenarios [36, 20]—to blend instances of these semantically related classes. This not only amplifies the presence of minority class samples but also encourages the model to learn finer distinctions between closely related categories (Figure 3).
In parallel, the RCM guides the formulation of our Semantic-Aware Loss, a weighted loss function that assigns greater importance to minority classes—especially those prone to misclassification. By amplifying the penalty for errors on these classes, our approach effectively mitigates the model’s inherent biases and promotes balanced, equitable learning across all categories.
4.1 Relation Contextual module
Previous efforts in addressing class imbalance within domain adaptation settings [50, 56, 104] have contributed significantly to improving performance on underrepresented classes. However, a common limitation across these methods is the neglect of inter-class relationships particularly how semantic similarity between classes influences misclassification patterns.
Our empirical analysis reveals that misclassifications of minority classes often occur in favor of semantically similar majority classes. To capture and leverage these subtle relationships, we introduce the Relation Contextual Module (RCM). Unlike general class imbalance metrics that can be inferred from data distributions, inter-class dynamics must be learned from the model’s behavior during training. RCM facilitates this by generating a confusion matrix at each training batch, comparing the model’s predictions with ground-truth labels. This matrix is normalized with respect to the ground truth, enabling a dynamic estimation of misclassification biases between specific class pairs. To ensure stability and robustness over time, we apply an Exponential Moving Average (EMA) to iteratively update a global confusion matrix. This evolving matrix serves as a reliable representation of the model’s long-term class bias patterns. Beyond providing smoother updates, the use of EMA eliminates the requirement for all classes to be present in every batch—thus streamlining the learning process. The procedure for updating this matrix is detailed in Algorithm 1.
When applied to target domain samples, RCM relies on high-confidence pseudo-labels to estimate class biases in the absence of ground-truth annotations, offering a reliable approximation that mirrors the supervised setup on the source domain.
4.2 Target Division and Augmentation
Aligning source and target domains is common in traditional domain adaptation tasks [58, 152, 116]. While alignment can be performed in either data space [12] or feature space [99], the absence of source data presents a unique challenge for domain alignment.
Even though the source data is inaccessible, the source-pretrained model retains crucial knowledge about the source domain. Following [21], we partition the target dataset into two groups, leveraging the pre-trained model to establish an explicit source-target domain distinction. To achieve this, a detection variance-based criterion is used, where variances are computed from the pre-trained model’s predictions on the target samples. A higher variance suggests a stronger resemblance to the source domain [21]. In particular, the model exhibits greater uncertainty (hard samples) when predicting source-similar images, leading to elevated variance values compared to source-dissimilar images. The detection variance is determined using the following formulation:
$v_{i}=\mathbb{E}[(F_{\theta_{s}}(X_{i})-\mathbb{E}[F_{\theta_{s}}(X_{i})])^{2}],$ (2)
where $F_{\theta_{s}}(X_{i})$ are the predictions of image $X_{i}$ via the source-pretrained model. Since this calculation is computationally intractable, we approximate it with Monte-Carlo sampling using dropout, following the method proposed by Gal and Ghahramani [35]. This approximation is achieved by conducting $M$ stochastic forward passes while keeping the detection model unchanged [9].
Since the outputs $F_{\theta_{s}}(X_{i})=(\mathbf{b_{i}},\mathbf{c_{i}})$ consist of bounding box coordinates and class scores, the detection variance is defined as the product of the bounding box variance $v_{bi}$ and the class score variance $v_{ci}$. For a given prediction with $N_{i}$ bounding boxes and $K$ classes, where $\mathbf{\{b_{ij}}=(x_{ij}^{1},y_{ij}^{1},x_{ij}^{2},y_{ij}^{2})\}_{j=1}^{N_{i}}$ and $\mathbf{\{c_{ij}}=(c_{ij}^{1},c_{ij}^{2},\ldots,c_{ij}^{K})\}_{j=1}^{N_{i}}$, we can express $v_{bi}$ and $v_{ci}$ as follows:
$v_{bi}=\frac{1}{MN_{i}}\sum_{j=1}^{N_{i}}\sum_{m=1}^{M}\|\mathbf{b_{ij}^{m}}-\mathbf{\bar{b}_{ij}}\|^{2},$ (3)
$v_{ci}=\frac{1}{MN_{i}}\sum_{j=1}^{N_{i}}\sum_{m=1}^{M}\|\mathbf{c_{ij}^{m}}-\mathbf{\bar{c}_{ij}}\|^{2},$ (4)
where $\mathbf{b_{ij}^{m}}$ and $\mathbf{c_{ij}^{m}}$ represent the localization coordinates and classification scores of the $m$-th forward pass for the $j$-th bounding box in $X_{i}$, respectively, and $\mathbf{\bar{b}_{ij}}$, $\mathbf{\bar{c}_{ij}}$ denote their corresponding average values over all $M$ forward passes.
The detection variance for an image $X_{i}$ is computed as $v_{i}=v_{bi}v_{ci}$. We then rank the images based on their variances, where $r_{i}$ represents the rank of $X_{i}$. The variance level $vl_{i}$ for the $i$-th image is given by $vl_{i}=\frac{r_{i}}{N}$. If $vl_{i}\geq\sigma$, we categorize $X_{i}$ as source-similar; otherwise, it is deemed source-dissimilar, with $\sigma\in(0,1)$ being a predefined threshold. This process effectively partitions the target domain into source-similar and source-dissimilar subsets for query-based adversarial alignment between domains.
Oversampling is a widely used strategy in classification tasks to mitigate class imbalance by increasing the representation of minority classes. However, this method presents challenges in object detection, where individual images often contain multiple object classes. Our analysis of breast cancer detection datasets such as DDSM [66] reveals that most images do not consistently feature even a single instance of majority classes, making conventional image-level resampling approaches largely ineffective. This calls for more targeted, instance-aware augmentation techniques tailored to object detection.
Following the instance-level augmentation strategy proposed by Zhang et al. [140], we extract object instances using bounding box annotations and strategically insert them into other images. To guide this process, we leverage our Relation Contextual Module (RCM) to identify majority and minority classes based on their likelihood of correct classification.
We compute the average diagonal value of the RCM —representing the mean self-classification probability—using the formula:
${\rm\texttt{RCM}}_{\rm avg}=\frac{1}{C}\sum^{C}_{c=0}{\texttt{RCM}}(c,c),$ (5)
where $C$ is the number of classes. Classes with ${\rm\texttt{RCM}}(c,c)$ values above this mean are labeled as majority classes, while those below are considered minority.
Unlike prior methods that randomly overlay instances, we employ a relation-guided MixUp strategy. For each base instance, we select a paired instance through weighted random sampling, where the weights are determined by inter-class misclassification probabilities captured in the RCM. For majority base instances, we use the column vector ${\rm\texttt{RCM}}(:,c)$, setting ${\texttt{RCM}}(c,c)=0$ to exclude self-augmentation and focus on classes often confused with the majority. For minority base instances, the row vector ${\rm\texttt{RCM}}(c,:)$ is used without masking, allowing for beneficial self-augmentation.
Paired instances are resized to match the base instance for bounding box consistency. MixUp is then applied as:
$\hat{I}$ $=\beta\cdot I_{\rm base}+(1-\beta)\cdot I_{\rm mix},$ (6)
$\hat{c}$ $=\beta\cdot c_{\rm base}+(1-\beta)\cdot c_{\rm mix},$
where $I_{\rm base}$ and $I_{\rm mix}$ are the cropped images of the base and mixed instances, and $\hat{I}$ is the resulting augmented image. Correspondingly, $c_{\rm base}$ and $c_{\rm mix}$ denote the respective class vectors, with $\hat{c}$ being the interpolated class label.
To support domain adaptation, we apply this augmentation to both source similar and dissimilar domains. For source-like images, we sample instances from both domains to promote knowledge transfer. For source dissimilar images, we prioritize dissimilar domain instances, falling back to source similar instances only when necessary. Importantly, we refrain from augmenting minority base instances in the target domain to preserve their semantic integrity, helping the model maintain focus on real target distributions and avoid domain drift.
Algorithm 1 Relation Contextual Module (RCM) Update Procedure
0:  Global class-relation matrix $\mathbf{R}\in\mathbb{R}^{C\times C}$ initialized to zero; EMA decay factor $\beta\in[0,1]$.
1:  while training do
2:     Initialize batch-specific relation matrix $\mathbf{R}_{b}\in\mathbb{R}^{C\times C}$ to zero.
3:     for each training example with ground-truth label $c_{i}$ and predicted label $x_{i}$ do
4:        $\mathbf{R}_{b}[c_{i},x_{i}]\leftarrow\mathbf{R}_{b}[c_{i},x_{i}]+1$
5:     end for
6:     for each class $c\in\{1,\dots,C\}$ do
7:        Normalize $\mathbf{R}_{b}[c,:]$ to sum to 1.
8:        Update $\mathbf{R}[c,:]\leftarrow\beta\cdot\mathbf{R}[c,:]+(1-\beta)\cdot\mathbf{R}_{b}[c,:]$
9:     end for
10:  end while
To enable inter-semantic Augmentation, we maintain a dynamic repository of class-specific instance crops, referred to as the Cropbank [140]. These instance crops are extracted from bounding box annotations in labeled source images as well as pseudo-labeled target images. To facilitate domain-specific augmentation, separate Cropbanks are maintained for source similar and source dissimilar subsets. Each Cropbank is updated in a first-in-first-out (FIFO) manner to ensure sample diversity and freshness. This strategy is particularly advantageous for the source dissimilar Cropbank, where early-stage pseudo-labels may be less reliable; replacing them over time helps improve augmentation quality as training progresses.
4.3 Semantic-Aware Loss
To further mitigate class bias, we introduce a weighted parameter into the classification loss, guided by the Relation Contextual Module (RCM) for foreground classes. This weighting scheme emphasizes classes that are frequently misclassified as dominant majority classes, thereby directing the model’s attention toward improving performance on underrepresented or confused categories.
To accentuate this focus, we apply a non-linear transformation to the RCM values as follows:
$w_{i}=\begin{cases}\sqrt{1-\text{{RCM}}(c_{i},x_{i})},&\text{if }c_{i}=x_{i}\\ \sqrt{\frac{\text{{RCM}}(c_{i},x_{i})}{\text{{RCM}}(c_{i},c_{i})}},&\text{otherwise}\end{cases}$ (7)
where $w_{i}$ denotes the weight for the $i^{\text{th}}$ instance, $c_{i}$ is its pseudo-label, and $x_{i}$ is the predicted class. When $c_{i}\neq x_{i}$, we normalize using the diagonal entry $\text{{RCM}}(c_{i},c_{i})$ to scale the misclassification likelihood relative to the class’s overall performance.
Background class weights are uniformly set to 1 to prevent model bias against them. To reconcile the disparity between foreground and background weights, we normalize foreground instance weights so that their mean equals that of the background class:
$W_{f}=\frac{W_{f}}{\text{mean}(W_{f})},$ (8)
where $W_{f}$ represents the set of weights for foreground instances.
To avoid excessive influence from extreme weights, we apply a regularization term $\lambda_{l}$ to all class-relation weights:
$W=\frac{W+\lambda_{l}}{1+\lambda_{l}},$ (9)
ensuring smoother gradients and preventing the loss from being overly sensitive to outlier weights.
The weighted classification loss is then defined as:
$\mathcal{L}_{\text{cls}}=\frac{1}{N}\sum_{i=1}^{N}w_{i}\cdot\text{{CE}}(c_{i},x_{i}),$ (10)
where $N$ is the total number of instances, and $\text{{CE}}(\cdot)$ denotes the standard cross-entropy loss.
4.4 Grounded Supervision
We introduce a framework that leverages Large Vision Foundation Models (LVFMs) through expert distillation, as depicted in 2. We introduce an Expert branch aimed at assimilating the strengths of LVFMs into a unified learning approach. As an initial assumption, we posit that the expert model can effectively represent a wide variety of images sourced from datasets like ImageNet1K or ImageNet21k [26], LAION-400M [103], or DataComp1B [34]. We have considered many prominent families of teacher models: CLIP [90], DINOV2 [87], ViT [32]and SAM [64], GroundingDINO [78], and BioMedParse [146]. These models have exhibited outstanding performance across a wide array of tasks, with CLIP demonstrating broad proficiency, DINOV2 excelling in downstream dense tasks such as semantic segmentation under linear probe, and SAM particularly shining in open-vocabulary segmentation tasks. The newly introduced Expert branch incorporates the pretrained LVFMs. In the context of knowledge distillation, utilizing a single expert’s embedding directly for distilling knowledge to the student model may not suffice. We found that relying on an expert’s embedding directly for distilling knowledge to the student model is insufficient for effective pseudo-label correction during adaptation. Hence, we propose using expert predictions and introducing a pseudo-supervised loss to guide the student model out from the local minima.
Our objective is to transfer the expertise of the expert model to the student model. The expert model $E(\cdot)$ is frozen, while the student model $S(\cdot)$ is updated via gradient descent during adaptation. To supervise the student using pseudo labels from the expert, we employ a combination of a consistency loss (mean teacher loss) and a bounding box regression loss.
Let $x_{i}$ be an input image. The expert produces pseudo-labels in the form of bounding box predictions $E_{\text{bbox}}^{i}=E(x_{i})$, which are used to supervise the student’s prediction $S_{\text{bbox}}^{i}=S(x_{i})$. We define the loss as:
$\mathcal{L}_{\text{expert}}=$ $\lambda_{\text{cls}}\cdot\mathcal{L}_{\text{cls}}(S_{\text{bbox}}^{i},E_{\text{bbox}}^{i})$
$+\lambda_{\text{reg}}\cdot\mathcal{L}_{\text{reg}}(S_{\text{bbox}}^{i},E_{\text{bbox}}^{i})$ (11)
where $\mathcal{L}_{\text{cls}}$ is the classification loss (semantic aware loss), and $\mathcal{L}_{\text{reg}}$ is the bounding box regression loss (e.g., smooth L1 or GIoU loss). The weights $\lambda_{\text{cls}}$ and $\lambda_{\text{reg}}$ control the contribution of each term.
This unified loss ensures that the student model learns both semantic consistency and localization accuracy from the expert model via pseudo-supervision.
4.5 Overall Objective and Training Strategy.
The final objective combines supervised, unsupervised, and discriminator losses:
$\mathcal{L}=\lambda_{u}\mathcal{L}_{\text{stu}}+\lambda_{d}\mathcal{L}_{\text{dis}}+\mathcal{L}_{\text{Expert}},$ (12)
where $\lambda_{u}$ and $\lambda_{d}$ control the contributions of the unsupervised and discriminator losses, respectively, and $\mathcal{L}_{\text{stu}},\ \text{and}\ \mathcal{L}_{\text{Expert}}$ have been formulated in Eq. (1), and (11) respectively. As the teaching process continues, an adequate amount of pseudo-boxes will be produced, and the influence of $\lambda_{u}$ should be lowered to avoid the encoder over-fitting to pseudo-labels. As a result, We decay $\lambda_{d}$ as adaptation continues. More details are discussed in the Supplementary material.
5 Experiments
We conduct extensive experiments to assess the effectiveness of our approach across various challenging medical imaging datasets, as well as standard UDA and SFOD benchmarks. In the UDA setting, both source and target domain data are accessible during training, facilitating adaptation. Conversely, in the SFOD scenario, adaptation is performed using only a source-trained model without access to source domain data. Additionally, we perform ablation studies employing diverse exchange strategies to validate the efficacy of our proposed methods. We further analyze the promising results obtained by our method through comprehensive visualizations and component-wise evaluations.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
UMT [137] CVPR’21 0.01 0.09 0.15 0.21 0.204 0.319
SFA [114] MM’21 0.03 0.13 0.23 0.35 0.241 0.338
H2FA [126] CVPR’22 0.13 0.32 0.45 0.52 0.575 0.312
AQT [53] IJCAI’22 0.11 0.37 0.44 0.57 0.680 0.349
AT [71] CVPR’22 0.19 0.38 0.65 0.75 0.721 0.512
DDSM D-Adapt [55] ICLR’22 0.11 0.29 0.45 0.59 0.731 0.414
to HT [29] CVPR’23 0.17 0.49 0.61 0.69 0.704 0.362
INBreast CLIPGAP [109] CVPR’23 0.15 0.55 0.61 0.75 0.712 0.445
ConfMIX [84] WACV’23 0.19 0.47 0.58 0.73 0.737 0.409
MRT [148] ICCV’23 0.16 0.54 0.64 0.72 0.789 0.489
D-MASTER [6] MICCAI’24 0.25 0.61 0.70 0.82 0.808 0.524
Mexforer [114] MM’21 0.02 0.03 0.03 0.03 0.060 0.090
IRG [27] CVPR’23 0.05 0.05 0.07 0.09 0.110 0.120
GT (Ours) 0.06 0.45 0.65 0.92 0.589 0.758
Table 1: Results of adaptation from DDSM to INBreast.
5.1 Datasets
Following datasets were used in our experiments.
  1. 1.
    DDSM [66] A publicly available BCD dataset comprising 2,620 full mammography images with 1162 malignancies. For full detail, please refer to Supplementary material.
  2. 2.
    INBreast [85] A smaller BCD dataset with 410 mammography images from 115 patients, including 87 malignancies.
  3. 3.
    RSNA [11] The original RSNA dataset consists of 54,706 screening mammograms, containing 1,000 malignancies from 8,000 patients. We curated a subset named RSNA-BSD1K, comprising 1,000 mammograms with 200 malignant cases, annotated at the bounding box level by expert radiologists.
  4. 4.
    Cityscapes [24] Cityscapes is collected from urban scenes containing 2,975 images for training and 500 images for validation.
  5. 5.
    Foggy Cityscapes [101]. This dataset is constructed by a fog synthesis algorithm from Cityscapes. We use Cityscapes as source domain and Foggy Cityscapes with highest fog density(0.02) as target domain. Similar to Cityscapes, this dataset integrates fog and depth information into street view images.
  6. 6.
    BDD100k [135] Created at UC Berkeley, it contains 100K annotated 720p images (1/frame/video) across 10 classes from diverse US driving scenes, supporting robust detection and adaptation studies.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
D-Adapt [55] ICLR’21 0.04 0.12 0.18 0.29 0.439 0.263
AT [73] CVPR’22 0.16 0.28 0.35 0.42 0.486 0.338
DDSM H2FA [127] CVPR’22 0.03 0.13 0.18 0.36 0.634 0.236
to MRT [147] ICCV’23 0.32 0.52 0.69 0.72 0.741 0.352
RSNA Mexforer [114] MM’21 0.24 0.31 0.39 0.39 0.336 0.287
IRG [27] CVPR’23 0.16 0.25 0.37 0.39 0.308 0.235
GT (Ours) 0.10 0.43 0.58 0.91 0.781 0.530
Table 2: Results of adaptation from DDSM to RSNA.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
D-Adapt [55] ICLR’21 0.00 0.06 0.09 0.10 0.381 0.362
AT [73] CVPR’22 0.01 0.08 0.10 0.15 0.385 0.311
RSNA H2FA [127] CVPR’22 0.02 0.08 0.10 0.12 0.483 0.315
to MRT [147] ICCV’23 0.03 0.09 0.12 0.17 0.739 0.587
INBreast Mexforer [114] MM’21 0.02 0.03 0.03 0.03 0.060 0.090
IRG [27] CVPR’23 0.05 0.05 0.07 0.09 0.110 0.120
GT (Ours) 0.01 0.28 0.49 0.90 0.638 0.605
Table 3: Results of adaptation from RSNA to INBreast.
Method Venue SF Person Rider Car Truck Bus Train Mcycle Bicycle mAP
DA-Faster [18] CVPR’18 29.2 40.4 43.4 19.7 38.3 28.5 23.7 32.7 32.0
EPM [49] ECCV’20 44.2 46.6 58.5 24.8 45.2 29.1 28.6 34.6 39.0
SSAL [86] NIPS’21 45.1 47.4 59.4 24.5 50.0 25.7 26.0 38.7 39.6
SFA [114] MM’21 46.5 48.6 62.6 25.1 46.2 29.4 28.3 44.0 41.3
UMT [28] CVPR’21 33.0 46.7 48.6 34.1 56.5 46.8 30.4 37.3 41.7
D-adapt [55] ICLR’21 40.8 47.1 57.5 33.5 46.9 41.4 33.6 43.0 43.0
TIA [144] CVPR’22 34.8 46.3 49.7 31.1 52.1 48.6 37.7 38.1 42.3
PT [44] ICML’22 40.2 48.8 63.4 30.7 51.8 30.6 35.4 44.5 42.7
MTTrans [136] ECCV’22 47.7 49.9 65.2 25.8 45.9 33.8 32.6 46.5 43.4
SIGMA [68] CVPR’22 44.0 43.9 60.3 31.6 50.4 51.5 31.7 40.6 44.2
O2net [41] MM’22 48.7 51.5 63.6 31.1 47.6 47.8 38.0 45.9 46.8
AQT [53] IJCAI’22 49.3 52.3 64.4 27.7 53.7 46.5 36.0 46.4 47.1
AT [71] CVPR’22 43.7 54.1 62.3 31.9 54.4 49.3 35.2 47.9 47.4
TDD [45] CVPR’22 50.7 53.7 68.2 35.1 53.0 45.1 38.9 49.1 49.2
CIGAR [79] CVPR’23 45.3 45.3 61.6 32.1 50.0 51.0 31.9 40.4 44.7
CSDA [38] ICCV’23 46.6 46.3 63.1 28.1 56.3 53.7 33.1 39.1 45.8
HT [30] CVPR’23 52.1 55.8 67.5 32.7 55.9 49.1 40.1 50.3 50.4
MRT [147] ICCV’23 52.8 51.7 68.7 35.9 58.1 54.5 41.0 47.1 51.2
SFOD [69] AAAI’21 21.7 44.0 40.4 32.6 11.8 25.3 34.5 34.3 30.6
SFOD-Mosaic [69] AAAI’21 25.5 44.5 40.7 33.2 22.2 28.4 34.1 39.0 33.5
HCL [52] NIPS’21 38.7 46.0 47.9 33.0 45.7 38.9 32.8 34.9 39.7
SOAP [124] IJIS’21 35.9 45.0 48.4 23.9 37.2 24.3 31.8 37.9 35.5
LODS [67] CVPR’22 34.0 45.7 48.8 27.3 39.7 19.6 33.2 37.8 35.8
AASFOD [21] AAAI’23 32.3 44.1 44.6 28.1 34.3 29.0 31.8 38.9 35.4
IRG [110] CVPR’23 37.4 45.2 51.9 24.4 39.6 25.2 31.5 41.6 37.1
PETS [77] ICCV’23 46.1 52.8 63.4 21.8 46.7 5.5 37.4 48.4 40.3
DACA [145] IJCV’24 44.7 31.2 60.1 53.1 53.9 0.0 27.8 45.9 39.9
LPLD [133] ECCV’24 39.7 49.1 56.6 29.6 46.3 26.4 36.1 43.6 40.9
SF-UT [43] ECCV’24 40.9 48.0 58.9 29.6 51.9 50.2 36.2 44.1 45.0
GT (Ours) 42.7 55.4 61.7 40.7 62.0 54.6 39.1 53.0 50.8
Oracle 66.3 61.1 80.8 45.6 68.8 52.0 49.1 54.9 59.8
Table 4: Results of adaptation from Normal Cityscapes to Foggy weather Cityscapes (CF). “SF” refers to source-free setting. “Oracle” refers to the models trained by using labels during training.
5.2 Implementation Details
Building upon existing research [51, 70, 147, 88, 21, 77], we validate our method across state-of-the-art SFOD and UDA benchmarks. We utilize the Faster R-CNN [95] architecture with VGG-16 and ResNet-101 as backbone networks. The model’s hyperparameters are set as follows: an Exponential Moving Average (EMA) decay rate of 0.9996, beta-distribution parameters of $[0.5,0.5]$, an adversarial loss weight ($\lambda_{d}$) of 0.1, an unsupervised loss weight ($\lambda_{u}$) of 1.0, and a regularization term ($\lambda_{l}$) of 1.0. A hard threshold ($\tau$) of 0.8 is applied for pseudo-labeling. We implement weak-strong augmentation on both source and target images. The training process involves initially training the student model for 20,000 iterations using labeled source data. Subsequently, the teacher model is initialized with the student’s parameters and updated through EMA at each iteration. Training continues for an additional 60,000 iterations incorporating both labeled source and unlabeled target data. Our framework is built upon the Detectron2 [121] platform. Experiments are conducted with a batch size of 8 source and 8 target images, distributed across 4 NVIDIA A100 GPUs. Further experimental setup details are provided in the Supplementary material.
5.3 Evaluation Metrics
For Natural Images, we report the AP for each class and mAP score, following previous works. For Medical Images, we use the Free-Response Receiver Operating Characteristic (FROC) curves for detection, F1-score, and AUC for reporting our classification results. The curves provide a graphical representation of sensitivity/recall values at different False Positives per Image (FPI). We follow related works in this area [91] and consider a prediction as true positive if the center of the predicted bounding box lies within the ground truth box.
5.4 Comparison with Current SOTA Methods
We evaluate the performance of our proposed Grounded Teacher (GT) method against other approaches on all three medical benchmarks and the natural benchmark mentioned earlier for generalizability. Since UDA and SFDOAD share similar task settings, we conducted comparisons with both. Table 1, Table 2 and Table 3 present the comparison results on medical image datasets. Table 4 presents comparison of the natural dataset. Our proposed GT consistently outperforms existing all SOTA methods, demonstrating generalizability and significant improvements across both natural and medical settings.
DDSM to INBreast. Adaptation from large to small-scale medical datasets with different modalities. Here, we consider DDSM [66] dataset as the source domain and INBreast [85] as the target domain. Results are presented in Table 1. Our proposed method demonstrates superior performance across various False positives per Image (FPI) values compared to existing methods as displayed.
RSNA to INBreast. Adaptation across medical datasets with different machine-acquisitions. This is vital for enhancing healthcare outcomes, improving diagnostic accuracy, and facilitating better clinical decisions. To evaluate our method’s performance, we adapt a model trained on DDSM [66] to RSNA [11]. Results for all FPI values are presented in Table 2, demonstrating that our method achieves state-of-the-art performance on this benchmark.
DDSM to RSNA. This experiment evaluates domain adaptation across datasets collected from distinct geographical regions. Specifically, we consider the DDSM [66] dataset as the source domain and RSNA [11] as the target domain. As shown in Table 2, our proposed method consistently outperforms existing approaches across all FPI thresholds, demonstrating its robustness and effectiveness in cross-domain generalization.
Cityscapes to Foggy Cityscapes. Object detectors often experience a significant drop in performance when deployed under adverse real-world conditions such as fog, due to the domain shift caused by the lack of such conditions in the training data. Domain adaptation aims to bridge this gap between normal and adverse weather scenarios. To investigate this, we conduct experiments on the widely-used Cityscapes $\rightarrow$ FoggyCityscapes benchmark. As presented in Table 4, student-teacher based frameworks consistently outperform non-student-teacher approaches by a notable margin. Our method, GT, built upon mean teacher architecture [71], achieves a substantial improvement with a performance of 50.8 mAP. Furthermore, it demonstrates notable gains in minority object classes, highlighting its robustness and adaptability in challenging weather conditions.
5.5 Ablation Studies
Method RCM SA SAL mAP
Base (AT [72]) 30.4
+ SA 41.8
+ SAL 44.2
GT (Full) 50.8
Table 5: Ablation studies on GT components. RCM is included in all studies as it forms the basis of CRA and SAL. We report the mean average precision at 0.50 IoU (mAP). Our contributions are not included in the base framework (AT [71]).
To verify the significance of our contributions, we conducted an ablation study. All experiments within this study were performed on the Cityscapes $\to$ Foggy Cityscapes benchmark using the VGG16 backbone.
5.5.1 Quantitative Ablation
Table 5 presents a quantitative analysis of the impact of each component within our framework. The base framework prior to integrating our modules aligns with the AT model [71], achieving a 30.4 mAP. Given that the Relation-Contextual Module (RCM) is integral to both class-relation augmentation and semantic-aware loss, it remains a constant across all experimental variations. Integrating Semantic Augmentation (SA) along with RCM increases the performance to 41.8 mAP, a gain of +11.4 over the base. Adding the Semantic Aware Loss (SAL) with RCM separately results in a higher mAP of 44.2, demonstrating a +13.8 improvement compared to the baseline, indicating that SAL provides a greater performance boost than SA in this ablation setting. Notably, Class-Relation Augmentation (CRA) significantly reduces the performance disparity between minority and majority classes, as evidenced. Finally, our full GT model, which combines RCM, SA, and SAL, achieves the highest score of 50.8 mAP, which not only boosts overall performance but also underscores our method’s efficacy in managing class imbalance.
Figure 3: FROC analysis of model generalization on unseen target domains. Performance comparison between our method (GT) and the MT baseline for detection tasks under domain shift: (a) DDSM→INBreast (410 images), (b) DDSM→RSNA (993 images), and (c) RSNA→INBreast (410 images), where our method demonstrates superior performance across different domain shifts. Higher curves indicate better trade-offs between sensitivity (TPR) and false positives per image (FPI).
Figure 4: Qualitative results of our proposed Grounded Teacher on Medical settings (DDSM $\rightarrow$ INBreast). We present visual comparisons between AT (top row) and GT (bottom row). Our method, GT, effectively identifies Malignancies better and mitigates false negatives. Bounding box colors indicate: Blue — Ground truth, Green — true positive, and Red — false negative.
Figure 5: Qualitative results of our proposed Grounded Teacher on Natural settings (Cityscapes $\rightarrow$ Foggy Cityscapes). We present visual comparisons between AT (top row) and GT (bottom row). Our method, GT, effectively mitigates misclassifications, reduces false negatives, and eliminates false positives. Bounding box colors indicate: Green — true positive, Blue — misclassified, Red — false negative, and Pink — false positive.
5.5.2 Impact of Augmentation
In our study, we assess the influence of augmentation strategies, focusing on both the proportion of augmented data and the criteria for selecting instances to augment. Image augmentation plays a pivotal role in our methodology, enhancing the dataset by incorporating additional representations, particularly of minority classes. However, excessive augmentation can lead to the model learning less accurate class representations. To mitigate this, we judiciously augment a random subset of images.
The pairing of class instances during augmentation is crucial for improving its effectiveness. Unlike random pairing, our Semantic Augmentation (SA) method prioritizes pairing instances from closely related minority and majority classes. This targeted approach ensures that the Mixup outputs are more meaningful, thereby enhancing the performance of minority classes.
Table 6 presents the experimental results of this approach. We compare the effects of randomly selecting class instances for Mixup against our SA method across different augmentation probabilities, which represent the likelihood of an instance in the base image being augmented. Applying Mixup randomly yields a modest improvement of +1.1 mAP in overall performance. In contrast, employing SA leads to a more substantial increase of +4.4 mAP. These findings indicate that pairing highly related classes during Mixup enhances minority class performance with minimal impact on majority classes. Furthermore, our experiments reveal that excessive augmentation can negatively affect model performance, underscoring the importance of a balanced augmentation strategy.
Selection Method mAP
Random (0.5) 46.4
Semantic Augmentation (1.0) 47.5
Semantic Augmentation (0.5) 50.8
Table 6: Comparison of selecting class instances randomly and via CRA. Values in brackets refer to the likelihood of an instance being augmented.
5.5.3 Weighing Strategy for Semantic Loss
To enhance the performance of minority classes, we introduce a Semantic Aware Loss (SAL) function, as defined in Eq 10. Class-level loss functions are a prevalent strategy for addressing dataset class imbalances. We compare our SAL with a variant of existing class-level losses that utilize only the diagonal elements of the Inter-Class Relation module, which correspond to the ground-truth class likelihood for accurate classification. In contrast, SAL leverages the likelihood of the ground-truth being classified as the predicted class to influence its loss function.
As shown in Table 7, incorporating inter-class relationships through SAL achieves an mAP of 50.8, a performance improvement of +3.7 mAP compared to the base. While SAL effectively utilizes inter-class information, there are instances where it might excessively penalize well-performing classes without proper constraints. To mitigate this, we apply a regularization term, as specified in Eq 9. Omitting this regularization (SAL w/o Reg.) results in an mAP of 47.8, demonstrating a significant performance drop of -3.0 mAP compared to the full SAL. This clearly highlights the importance of the regularization term, likely preventing certain class weights from becoming disproportionately small during training.
Class Weight Strategy mAP
Class-Level 47.1
SAL w/o Reg. 47.8
SAL 50.8
Table 7: Class loss weighting strategies. Class-Level uses only the diagonal values in our RCM with regularization. SAL refers to our full Semantic Aware Loss.
5.5.4 Qualitative Results
Figure 5 presents the qualitative outcomes of our method, comparing predictions from baseline AT [71] in the top row with those from the Ground Truth (GT) in the bottom row. The GT effectively corrects misclassifications, as highlighted by the blue boxes. For instance, in columns 1, 2, and 4, the GT accurately labels ’person’, ’car’, and ’truck’, respectively, rectifying the AT’s errors. Moreover, the GT demonstrates a notable reduction in both false positives and false negatives, indicated by the pink and red boxes, respectively. This enhancement signifies improved detection accuracy across various object scales and classes. This improvement is particularly evident in column 3, where the GT successfully identifies small-scale objects and provides more accurate detections for larger-scale objects. Also in Fig 5, we show how qualitatively the predictions are improved with various proposed modules.
6 Limitations
We also evaluated the Cityscapes → BDD100k setup to test further on diverse domains. However, we observed subpar performance on this setup due to several inherent challenges. First, BDD100k is significantly larger with 100k annotated images labeled object instances across 10 categories [135]. In contrast, Cityscapes contains less than 3k training images with fewer 8 classes and more consistent urban scenes. The increased scale, greater intra-class diversity, and higher number of object categories in BDD100k likely introduce a domain gap too wide for effective unsupervised adaptation under our current framework. Our architecture, when trained on Cityscapes and tested directly on BDD100k does not surpass existing state-of-the-art (SOTA) SFOD methods, but it does show improvement with adaptation over source-only baselines as noted in Table 8.
Table 8: Comparison of results of adaptation from Cityscapes to BDD100k (CB) with source-free setting baselines
Method mAP
Source-trained 20.2
SFOD [69] 26.6
SFOD-Mosaic [69] 29.0
HCL [52] 30.3
GT (Ours) 29.1
Considering our motivation and primary contribution lie in adapting models for medical image analysis, our method achieves promising results across multiple breast cancer detection datasets, validating its effectiveness in the domain. The challenges observed in the Cityscapes → BDD100k setup while the simultaneous improvement with adaptation over source-only baselines highlight opportunities for future work regarding which, based on our findings, we believe that further investigation into inter-class dynamics is a promising direction for advancing context imbalance in the SFOD setting as a whole, especially in medical scenarios.
7 Conclusion
We have proposed a novel source-free domain adaptive object detection framework that leverages a Relation Contextual module and Expert Branch, effectively approximating and mitigating class biases, leading to more equitable performance across classes. Through comprehensive experiments conducted across various natural and medical image benchmark datasets, we consistently demonstrate that our technique outperforms the state-of-the-art. Our ablation studies confirm the effectiveness of each design aspect in improving the model’s domain adaptation capability.
Data Availability Statement. All results presented in this paper are based on publicly available datasets. To ensure reproducibility, we are making the processed data openly available. The datasets, along with the associated processing scripts, will be publicly released at the following GitHub repository at this link.. This repository will assist other researchers in replicating the results and conducting further experiments.
Appendix
Further Details on Datasets.
Our experiments are conducted on six datasets. Along with this, we introduce the SFOD problem in the medical domain and conduct experiments on three Breast Cancer Detection (BCD) datasets.
1) Cityscapes [23] is gathered from urban environments across 50 European cities, provides detailed annotations for 30 semantic classes across 8 categories. It comprises 5,000 high-quality annotated images and a larger set of 20,000 coarsely annotated images, all high-resolution (2048x1024 pixels) providing detailed visual data for precise scene understanding. Our study utilizes the high-quality subset of 5,000 images, consisting of 2,975 training images and a standard test set of 500 images from Frankfurt, Munster, and Lindau, as employed in prior research. This dataset is release under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License, and is available for academic and research purposes.
2) Foggy Cityscapes [102] is an extension of the Cityscapes dataset, designed to support research in developing robust computer vision algorithms for autonomous driving in foggy conditions. Similarly, it consists of high-resolution images (2048x1024 pixels) with annotations ofr 2D bounding boxes, pixel-level semantic segmentation, and instance segmentation, inherited from the original dataset. It is directly constructed from Cityscapes by simulating three levels of foggy weather (0.005,0.001,0.02), but we adapt on the most extreme level (0.02) in our experiments as done by Kennerley et al. [60]. This dataset is released under the Creative Commons Attribution-NonCommercial-ShareAlike 4.0 International License.
3) INBreast [85] is a relatively small breast cancer detection dataset, consisting of 410 mammography images from 115 patients, including 87 confirmed malignancies. The training images include both histologically confirmed cancers and benign lesions initially recalled for further examination but later identified as nonmalignant. We hypothesize that incorporating both malignant and benign lesions in the training process will enhance our model’s ability to detect a broader range of lesions and effectively distinguish between malignant and benign cases.
Table 9: Details of Training Datasets
Dataset Type Pre-training (Source) Unsupervised Adaptation (Target)
Train Val/Test Train Test
Cityscapes Natural 2,975 500 2,975 500
Foggy Cityscapes Natural - - 2,975 500
BDD100k Natural - - 36,728 5,258
RSNA Medical 895 (170) 98 (18) 993 (188) 993(188)
INBreast Medical - - 410 (91) 410 (91)
DDSM Medical 2885 (1339) 218 (118) 3103 (1458) 3103 (1458)
4) DDSM [66] is a publicly available breast cancer detection dataset, comprising 2,620 full mammography images, with 1,162 containing malignancies. The DDSM dataset offers digitized film-screen mammography exams with lesion annotations at the pixel level, where cancerous lesions are histologically confirmed. We used the DDSM dataset exclusively for training our model and not for evaluation. This decision stems from the observation that the quality of digitized film-screen mammograms is inferior to that of full-field digital mammograms, making evaluation on these cases less relevant. For our purposes, we converted the lossless JPEG images to PNG format, mapped pixel values to optical density using calibration functions from the DDSM website, and rescaled the pixel values to a range of 0–255.
5) RSNA-BSD1K The RSNA dataset [11] was contributed by mammography screening programs in Australia and the U.S. It contains radiographic breast images of female subjects in dicom format. It includes detailed labels, with radiologists’ evaluations and follow-up pathology results for suspected malignancies. The original RSNA dataset consists of 54,706 screening mammograms, containing 1,000 malignancies from 8,000 patients. We curated a subset named RSNA-BSD1K, comprising 1,000 mammograms with 200 malignant cases, annotated at the bounding box level by 2 expert radiologists.
5) BDD100k The BDD100k dataset [135], created by researchers primarily at UC Berkeley to address the limitations of prior driving datasets in scale and diversity, provides 100,000 diverse, 720p driving videos (split 70K train/10K val/20K test) from varied US locations, weather, and times. For object detection, one frame per video (at the 10th second) is annotated with bounding boxes for 10 categories, totaling 100K annotated images. It features over 1.8 million annotated instances across categories like car (1M+), sign (343K+), light (265K+), and person (129K+). Visibility attributes indicate about 48% of instances are occluded and 7% truncated. Its scale (70K training images) and diversity support robust detector training and domain adaptation studies, and significantly boost performance in related tasks like instance segmentation and tracking when used in joint training.
It is important to note that unlike single domain detection techniques for natural images, where a subset of the target dataset is reserved for adaptation and another for testing, our approach operates without requiring any labeled samples from the target domain. This characteristic makes it particularly suitable for medical datasets, where it is more practical to utilize the entire dataset during both training and testing. Therefore, when we refer to training on “Dataset A to Dataset B,” we imply that the model is trained using the full source dataset ($\mathcal{D}_{s}=A$) and adapted to the full target dataset ($\mathcal{D}_{t}=B$) in a fully unsupervised manner. Table 9 outlines the specific dataset splits used in our experiments.
Table 10: Data Augmentation Methods and Parameters
Augmentation Type Method Parameters
Weak Random Horizontal Flip Probability = 0.5
Resize Size = 800, Max Size = 1333
Strong Color Jitter (Color Adjustment) Brightness = 0.4
Contrast = 0.4
Saturation = 0.4
Hue = 0.1
Random Grayscale Probability = 0.2
Gaussian Blur Sigma Range = [0.1, 2.0
Probability = 0.5
Normalization Mean = [0.485, 0.456, 0.406]
Std = [0.229, 0.224, 0.225]
More Details on Augmentations
In our study, we employed both weak and strong augmentation techniques to enhance the robustness and generalization capabilities of our model. These augmentations are applied to the training images to simulate various real-world scenarios and improve the model’s performance.
Weak Augmentation. Weak augmentations are relatively simple transformations that slightly alter the images without significantly changing their content. We utilized two primary weak augmentation techniques. The first is Random Horizontal Flip, which flips the image horizontally with a probability of 0.5. This helps the model learn invariance to the left-right orientation of objects, making it more robust to such variations. The second technique is Resize, where images are resized such that the shortest edge is 800 pixels while maintaining the aspect ratio, and if the longest edge exceeds 1333 pixels, the image is scaled down accordingly. This resizing standardizes the input image size, ensuring consistent and efficient training. Together, these weak augmentations provide a baseline level of variability in the training data, helping the model to generalize better across different image scales and orientations.
Strong Augmentation. Strong augmentations involve more complex transformations that significantly alter the images, thereby providing a broader range of variability. These augmentations are designed to challenge the model and improve its ability to handle diverse and complex real-world scenarios. The first strong augmentation is Color Jitter, which randomly changes the brightness, contrast, saturation, and hue of the images with specified parameters (brightness = 0.4, contrast = 0.4, saturation = 0.4, hue = 0.1) and a probability of 0.8. This helps the model become invariant to different lighting conditions and color variations. The second augmentation is Random Grayscale, which converts the image to grayscale with a probability of 0.2, encouraging the model to focus on shapes and structures rather than colors. The third strong augmentation is Gaussian Blur, applied with a sigma range of [0.1, 2.0] and a probability of 0.5, simulating out-of-focus conditions and reducing high-frequency noise. Finally, all images are converted to tensors and normalized using the mean [0.485, 0.456, 0.406] and standard deviation [0.229, 0.224, 0.225]. These strong augmentations introduce substantial variability in the training data, forcing the model to learn more robust and generalized features.
Details on Hyperparameters
Table 11: Detailed hyper-parameters corresponding to each benchmark for Medical settings (DDSM $\rightarrow$ INBreast) and Natural settings (Cityscapes $\rightarrow$ Foggy Cityscapes)
Hyper-parameter Description DDSM $\rightarrow$ INBreast Cityscapes $\rightarrow$ Foggy Cityscapes - Detector FRCNN FRCNN - Backbone VGG VGG - BatchNorm True True - Expert BioMedParse [146] GroundingDINO [78] $N_{c}$ Number of categories for classification head 1 8 $\alpha$ Decay rate for student-teacher EMA 0.9996 0.9996 $\beta$ Beta-distribution parameters for mixup [0.5,0.5] [0.5,0.5] $\lambda_{d}$ Weight for Adversarial Loss 0.1 0.1 $\lambda_{u}$ Weight for Student Loss 1.0 1.0 $\lambda_{e}$ Weight for Expert Loss 1.0 1.0 $\tau$ Threshold value for pseudo-label confidence 0.8 0.8 $\lambda_{I}$ Regularization term for Semantic Aware Loss SAL 1.0 1.0 - Source-similar Augmentation Ratio 0.5 0.5 - Source-dissimilar Augmentation Ratio 0.5 0.5 - Burn-Up Step Iterations 20000 20000 - Total Training Iterations 40000 80000 $base\_lr$ Learning Rate 0.01 0.02
The hyperparameters used for training the models on the Medical setting (DDSM $\rightarrow$ INBreast) and Natural setting (Cityscapes $\rightarrow$ Foggy Cityscapes) domain adaptation task are summarized in Table 11. We employ the Adam optimizer with a base learning rate of $1\times 10^{-4}$ and a weight decay of $1\times 10^{-5}$ to ensure stable training. The models are trained for 50 epochs with a batch size of 16, and the input image resolution is fixed to $512\times 512$ pixels.
To balance the contributions from different training objectives, we set the unsupervised loss weight $\lambda_{u}$ to 2.0 and the discriminator loss weight $\lambda_{d}$ to 0.5. The expert loss is composed of a classification and regression component, both weighted equally with $\lambda_{\text{cls}}=1.0$ and $\lambda_{\text{reg}}=1.0$. For data augmentation, we apply standard techniques including random flipping, scaling, and color jittering to improve generalization.
These hyperparameters were found to be effective for domain adaptation under adverse weather conditions, and we retain them consistently across all comparative experiments unless explicitly varied in ablation studies.
Additional Visualizations
(a) Aachen
(b) Zurich
(c) Strasbourg
Figure 6: Qualitative results comparing our proposed Grounded Teacher (GT) with the baseline (AT) on Natural settings (Cityscapes $\rightarrow$ Foggy Cityscapes) using challenging foggy driving scenes from (a) Aachen, (b) Zurich, and (c) Strasbourg. Each sub-figure presents the AT output (top row) and the corresponding GT output (bottom row). Our method (GT) demonstrates improved performance by mitigating misclassifications, reducing false negatives, and eliminating false positives compared to AT. Bounding box colors indicate detection status: Green — true positive, Blue — misclassified, Red — false negative, Pink — false positive.
(a) Cologne
(b) Krefeld
(c) Monchengladbach
Figure 7: (Contd.) Qualitative results comparing our proposed Grounded Teacher (GT) with the baseline (AT) on Natural settings (Cityscapes $\rightarrow$ Foggy Cityscapes) using challenging foggy driving scenes from (a) Cologne, (b) Krefeld, and (c) Monchengladbach. Each sub-figure presents the AT output (top row) and the corresponding GT output (bottom row). Our method (GT) demonstrates improved performance by mitigating misclassifications, reducing false negatives, and eliminating false positives compared to AT. Bounding box colors indicate detection status: Green — true positive, Blue — misclassified, Red — false negative, Pink — false positive.
(a) DDSM $\rightarrow$ INBreast
(b) DDSM $\rightarrow$ RSNA
(c) RSNA $\rightarrow$ INBreast
Figure 8: Qualitative results comparing our proposed Grounded Teacher (GT) with the baseline (AT) on Medical settings using challenging Mammograms adapted from (a) DDSM $\rightarrow$ INBreast, (b) DDSM $\rightarrow$ RSNA, and (c) RSNA $\rightarrow$ INBreast. Each sub-figure presents the AT output (top row) and the corresponding GT output (bottom row). Our method (GT) demonstrates improved performance by effective identification of Malignancies and mitigating false negatives compared to AT. Bounding box colors indicate detection status: Blue — Ground truth, Green — true positive, Red — false negative.