### v2

Abstract
[…]
keywords:
Generalization, Source-free Domain Adaptation, Large Vision Foundation Models, Object Detection
1 Introduction
Unsupervised Domain Adaptation (UDA) offers a strategy for adapting object detectors to new domains where labeled data is not available.
Moreover many popular SFOD techniques utilize a self-supervised approach in a student-teacher (ST) framework. These methods bootstrap by training on pseudo-labels generated by a source-pretrained model  [67, 123, 70, 21, 112]. However, if the source data is biased, or the domain shift between source and target domains is significant, there is noise in the pseudo-labels, which impacts the training of a student model [27]. Since the pseudo-label error is significant, the, Exponential Moving Average (EMA) step, which updates the teacher model from the student model’s weight, ends up corrupting the teacher model as well leading to a mode collapse [3]. This phenomenon of progressive degradation, especially in source-free setting where source data is not available during adaptation, is visually evident in Figure 1, where traditional Mean Teacher frameworks and SF-Student-Teacher frameworks suffer from declining detection quality over time, while our method remains robust over the epochs.
Figure 1: Effect of pseudo-label noise and EMA updates on predictions across epochs for different SFOD methods (ST, SF-ST, and Ours). ST refers to Mean Teacher [105] and SF-ST refers to source-free setting. Both exhibit progressive degradation, while our proposed GT method preserves accurate and consistent performance over time.
To address these challenges, in this paper, we introduce our Grounded Teacher (GT), specifically designed to tackle context imbalance and mode collapse in the SFOD setting. Our approach introduces a Relation-Contextual Module (RCM) that explicitly models class biases and inter-class dynamics within the training data. Leveraging these insights, we implement a semantic augmentation strategy that enhances minority class representation by strategically blending them with visually similar majority class instances, drawn from our Cropbank repository[140]. This augmentation operates bidirectionally across domains, simultaneously addressing both class imbalance and domain shift. To further mitigate inter-class bias, we propose a Semantic-Aware Loss that dynamically prioritizes minority classes during training, particularly focusing on challenging cases where they are most vulnerable to misclassification as majority classes. Together, these components form a comprehensive framework that holistically addresses representation learning challenges in cross-domain scenarios.
Another key insight of this paper is that one can use learnt representations from Vision Foundational Models to guide a student model in SFOD settings. Due to the excellent zero shot capabilities of the Vision Foundation Models (VFMs), they can effectively guide student models to learn the right representation, and distances between the samples, even when the target domain has a large shift from the source domain. This helps the model overcome mode collapse caused by biased pseudo-labels during unsupervised training.
Through the integration of these approaches, we observe enhanced pseudo-label quality, leading to measurable gains on medical imaging benchmarks achieving a new state-of-the-art (SOTA) of 50.8 mAP in natural images for Cityscapes $\to$ Foggy Cityscapes, beating the previous standard of 45.0 mAP by +5.8 mAP ($\sim$13performance gain). We summarize the contributions of this paper as follows:
  • We highlight the problem of training instability of a student model in SFOD setting due to incorrect pseudo labels from a source pretrained model in the presence of source bias, or large domain shift between source and target domains.
  • We propose our GT model, supported by our RCM, which can map the model’s existing class biases. To this end, we formulate a Semantic Aware Loss (SAL) to further enhance the performance of biased classes.
  • We propose to leverage the zero-shot performance of VFMs to guide the student model in a sample-specific manner and help learn the correct representation in a large domain shift scenario.
2 Preliminaries
Problem statement. Let $X$ denotes the input space and $\mathcal{D}_{s}=\{(x_{s}^{i},y_{s}^{i})\}_{i=1}^{N_{s}}$, the source domain, where $y_{s}^{i}=(\mathbf{b_{s}^{i}},c_{s}^{i})$ is a tuple comprising of ground truth bounding box ($\mathbf{b_{s}^{i}}\in\mathbf{\mathcal{R}^{4}}$) and the corresponding class label ($c_{s}^{i}\in\mathbf{\mathcal{R}}$) of the object respectively. Here, $x_{s}\in X$ denotes the image from the source domain, and $N_{s}$ denotes the total number of source images. The target domain $\mathcal{D}_{t}=\{x_{t}^{i}\}_{i=1}^{N_{t}}$ is unlabeled, where $N_{t}$ denotes the total number of target images, and the target samples $x_{t}^{i}\in X$. Our objective is to transfer the source-trained model to the target domain without accessing samples from the source domain.
Mean Teacher Framework. The self-supervised adaptation strategy involves updating the student model using unlabeled target data, leveraging pseudo-labels generated from the teacher model. Both the student and the teacher models are initialized by the source-trained model. The pseudo labels undergo a confidence threshold-based filtering process, and only the reliable ones are utilized to supervise the student training [62]. The pseudo-label supervision loss for the object detection model can be expressed as:
$\mathcal{L}_{\text{stu}}=\mathcal{L}_{\text{box}}^{S}(x_{t}^{i},yt_{t}^{i})+\mathcal{L}_{\text{giou}}^{S}(x_{t}^{i},yt_{t}^{i})+\mathcal{L}_{\text{cls}}^{S}(x_{t}^{i},yt_{t}^{i}).$ (1)
Here $yt_{t}^{i}$ represents the pseudo-label for image $i$ obtained after filtering low-confidence predictions from the teacher. The teacher is updated progressively through an Exponential Moving Average (EMA) of student weights. We implement a hard threshold, $\tau$, on the classification scores generated by the teacher to ensure that only pseudo-labels with high confidence are utilized by the student network, thereby encouraging more reliable learning outcomes.
Thus, the overall self-training process for the traditional student-teacher (ST) based object detection framework is expressed as follows:
$\Theta_{s}$ $\leftarrow\Theta_{s}+\gamma\frac{\partial L_{\text{stu}}}{\partial\Theta_{s}},$ (2)
$\Theta_{t}$ $\leftarrow\alpha\Theta_{t}+(1-\alpha)\Theta_{s}.$ (3)
Here $\mathcal{L}_{\text{stu}}$ represents the student loss computed using pseudo-labels from the teacher network, and $\Theta_{s}$ and $\Theta_{t}$ denote student and teacher network, respectively. The parameters $\gamma$ and $\alpha$ denote the student learning rate and teacher EMA rate, respectively. Although the ST framework enables knowledge distillation using noisy pseudo-labels, it alone is insufficient for learning high-quality target features in a source-free setting and leads to model degradation. Therefore, to enhance the features in the target domain, we introduce Visual Foundation Models (VFMs) as Expert and incorporate supervised loss for knowledge transfer from expert VFMs. Following [72], a discriminator is added to encourage domain invariant feature representations with an associated loss, $\mathcal{L}_{\rm dis}$.
Figure 2: Architecture of proposed Grounded Teacher (GT) Framework
3 Methodology
Our proposed method, Grounded Teacher (GT), illustrated in Figure 3, extends the standard mean-teacher framework by introducing a novel Relation Contextual Module (RCM). At the heart of GT, the RCM is designed to systematically capture and quantify class-specific biases in the model. Unlike conventional methods that address class imbalance in a general way, RCM explicitly models the semantic relationships between classes, with a particular focus on minority classes that are frequently misclassified as dominant ones.
This is achieved through the construction of a batch-wise confusion matrix based on pseudo-labels, which is normalized and continuously aggregated into a global matrix. This evolving matrix provides a robust and dynamic representation of the model’s bias landscape, forming the foundation for two key components of our method: Class-Relation Augmentation and Semantic Loss. Class-Relation Augmentation targets image-level class imbalance. Using the class relationships encoded by RCM, we identify minority classes in an image that share strong semantic overlap with dominant classes. To increase their representation and improve discrimination, we apply MixUp [142]—a technique known to be effective in imbalanced scenarios [36, 20]—to blend instances of these semantically related classes. This not only amplifies the presence of minority class samples but also encourages the model to learn finer distinctions between closely related categories (Figure 3).
In parallel, the RCM guides the formulation of our Semantic-Aware Loss, a weighted loss function that assigns greater importance to minority classes—especially those prone to misclassification. By amplifying the penalty for errors on these classes, our approach effectively mitigates the model’s inherent biases and promotes balanced, equitable learning across all categories.
3.1 Relation Contextual module
Previous efforts in addressing class imbalance within domain adaptation settings [50, 56, 104] have contributed significantly to improving performance on underrepresented classes. However, a common limitation across these methods is the neglect of inter-class relationships particularly how semantic similarity between classes influences misclassification patterns.
Our empirical analysis reveals that misclassifications of minority classes often occur in favor of semantically similar majority classes. To capture and leverage these subtle relationships, we introduce the Relation Contextual Module (RCM). Unlike general class imbalance metrics that can be inferred from data distributions, inter-class dynamics must be learned from the model’s behavior during training. RCM facilitates this by generating a confusion matrix at each training batch, comparing the model’s predictions with ground-truth labels. This matrix is normalized with respect to the ground truth, enabling a dynamic estimation of misclassification biases between specific class pairs. To ensure stability and robustness over time, we apply an Exponential Moving Average (EMA) to iteratively update a global confusion matrix. This evolving matrix serves as a reliable representation of the model’s long-term class bias patterns. Beyond providing smoother updates, the use of EMA eliminates the requirement for all classes to be present in every batch—thus streamlining the learning process. The procedure for updating this matrix is detailed in Algorithm 1.
When applied to target domain samples, RCM relies on high-confidence pseudo-labels to estimate class biases in the absence of ground-truth annotations, offering a reliable approximation that mirrors the supervised setup on the source domain.
Aligning source and target domains is common in traditional domain adaptation tasks [58, 152, 116]. While alignment can be performed in either data space [12] or feature space [99], the absence of source data presents a unique challenge for domain alignment.
Even though the source data is inaccessible, the source-pretrained model retains crucial knowledge about the source domain. Following [21], we partition the target dataset into two groups, leveraging the pre-trained model to establish an explicit source-target domain distinction. To achieve this, a detection variance-based criterion is used, where variances are computed from the pre-trained model’s predictions on the target samples. A higher variance suggests a stronger resemblance to the source domain [21]. In particular, the model exhibits greater uncertainty (hard samples) when predicting source-similar images, leading to elevated variance values compared to source-dissimilar images. The detection variance is determined using the following formulation:
$v_{i}=\mathbb{E}[(F_{\theta_{s}}(X_{i})-\mathbb{E}[F_{\theta_{s}}(X_{i})])^{2}],$ (2)
where $F_{\theta_{s}}(X_{i})$ are the predictions of image $X_{i}$ via the source-pretrained model. Since this calculation is computationally intractable, we approximate it with Monte-Carlo sampling using dropout, following the method proposed by Gal and Ghahramani [35]. This approximation is achieved by conducting $M$ stochastic forward passes while keeping the detection model unchanged [9].
Since the outputs $F_{\theta_{s}}(X_{i})=(\mathbf{b_{i}},\mathbf{c_{i}})$ consist of bounding box coordinates and class scores, the detection variance is defined as the product of the bounding box variance $v_{bi}$ and the class score variance $v_{ci}$. For a given prediction with $N_{i}$ bounding boxes and $K$ classes, where $\mathbf{\{b_{ij}}=(x_{ij}^{1},y_{ij}^{1},x_{ij}^{2},y_{ij}^{2})\}_{j=1}^{N_{i}}$ and $\mathbf{\{c_{ij}}=(c_{ij}^{1},c_{ij}^{2},\ldots,c_{ij}^{K})\}_{j=1}^{N_{i}}$, we can express $v_{bi}$ and $v_{ci}$ as follows:
$v_{bi}=\frac{1}{MN_{i}}\sum_{j=1}^{N_{i}}\sum_{m=1}^{M}\|\mathbf{b_{ij}^{m}}-\mathbf{\bar{b}_{ij}}\|^{2},$ (3)
$v_{ci}=\frac{1}{MN_{i}}\sum_{j=1}^{N_{i}}\sum_{m=1}^{M}\|\mathbf{c_{ij}^{m}}-\mathbf{\bar{c}_{ij}}\|^{2},$ (4)
where $\mathbf{b_{ij}^{m}}$ and $\mathbf{c_{ij}^{m}}$ represent the localization coordinates and classification scores of the $m$-th forward pass for the $j$-th bounding box in $X_{i}$, respectively, and $\mathbf{\bar{b}_{ij}}$, $\mathbf{\bar{c}_{ij}}$ denote their corresponding average values over all $M$ forward passes.
The detection variance for an image $X_{i}$ is computed as $v_{i}=v_{bi}v_{ci}$. We then rank the images based on their variances, where $r_{i}$ represents the rank of $X_{i}$. The variance level $vl_{i}$ for the $i$-th image is given by $vl_{i}=\frac{r_{i}}{N}$. If $vl_{i}\geq\sigma$, we categorize $X_{i}$ as source-similar; otherwise, it is deemed source-dissimilar, with $\sigma\in(0,1)$ being a predefined threshold. This process effectively partitions the target domain into source-similar and source-dissimilar subsets for query-based adversarial alignment between domains.
Oversampling is a widely used strategy in classification tasks to mitigate class imbalance by increasing the representation of minority classes. However, this method presents challenges in object detection, where individual images often contain multiple object classes. Our analysis of breast cancer detection datasets such as DDSM [66] reveals that most images do not consistently feature even a single instance of majority classes, making conventional image-level resampling approaches largely ineffective. This calls for more targeted, instance-aware augmentation techniques tailored to object detection.
Following the instance-level augmentation strategy proposed by Zhang et al. [140], we extract object instances using bounding box annotations and strategically insert them into other images. To guide this process, we leverage our Relation Contextual Module (RCM) to identify majority and minority classes based on their likelihood of correct classification.
We compute the average diagonal value of the RCM —representing the mean self-classification probability—using the formula:
${\rm\texttt{RCM}}_{\rm avg}=\frac{1}{C}\sum^{C}_{c=0}{\texttt{RCM}}(c,c),$ (5)
where $C$ is the number of classes. Classes with ${\rm\texttt{RCM}}(c,c)$ values above this mean are labeled as majority classes, while those below are considered minority.
Unlike prior methods that randomly overlay instances, we employ a relation-guided MixUp strategy. For each base instance, we select a paired instance through weighted random sampling, where the weights are determined by inter-class misclassification probabilities captured in the RCM. For majority base instances, we use the column vector ${\rm\texttt{RCM}}(:,c)$, setting ${\texttt{RCM}}(c,c)=0$ to exclude self-augmentation and focus on classes often confused with the majority. For minority base instances, the row vector ${\rm\texttt{RCM}}(c,:)$ is used without masking, allowing for beneficial self-augmentation.
Paired instances are resized to match the base instance for bounding box consistency. MixUp is then applied as:
$\hat{I}$ $=\beta\cdot I_{\rm base}+(1-\beta)\cdot I_{\rm mix},$ (6)
$\hat{c}$ $=\beta\cdot c_{\rm base}+(1-\beta)\cdot c_{\rm mix},$
where $I_{\rm base}$ and $I_{\rm mix}$ are the cropped images of the base and mixed instances, and $\hat{I}$ is the resulting augmented image. Correspondingly, $c_{\rm base}$ and $c_{\rm mix}$ denote the respective class vectors, with $\hat{c}$ being the interpolated class label.
To support domain adaptation, we apply this augmentation to both source similar and dissimilar domains. For source-like images, we sample instances from both domains to promote knowledge transfer. For source dissimilar images, we prioritize dissimilar domain instances, falling back to source similar instances only when necessary. Importantly, we refrain from augmenting minority base instances in the target domain to preserve their semantic integrity, helping the model maintain focus on real target distributions and avoid domain drift.
Algorithm 1 Relation Contextual Module (RCM) Update Procedure
0:  Global class-relation matrix $\mathbf{R}\in\mathbb{R}^{C\times C}$ initialized to zero; EMA decay factor $\beta\in[0,1]$.
1:  while training do
2:     Initialize batch-specific relation matrix $\mathbf{R}_{b}\in\mathbb{R}^{C\times C}$ to zero.
3:     for each training example with ground-truth label $c_{i}$ and predicted label $x_{i}$ do
4:        $\mathbf{R}_{b}[c_{i},x_{i}]\leftarrow\mathbf{R}_{b}[c_{i},x_{i}]+1$
5:     end for
6:     for each class $c\in\{1,\dots,C\}$ do
7:        Normalize $\mathbf{R}_{b}[c,:]$ to sum to 1.
8:        Update $\mathbf{R}[c,:]\leftarrow\beta\cdot\mathbf{R}[c,:]+(1-\beta)\cdot\mathbf{R}_{b}[c,:]$
9:     end for
10:  end while
To enable inter-semantic Augmentation, we maintain a dynamic repository of class-specific instance crops, referred to as the Cropbank [140]. These instance crops are extracted from bounding box annotations in labeled source images as well as pseudo-labeled target images. To facilitate domain-specific augmentation, separate Cropbanks are maintained for source similar and source dissimilar subsets. Each Cropbank is updated in a first-in-first-out (FIFO) manner to ensure sample diversity and freshness. This strategy is particularly advantageous for the source dissimilar Cropbank, where early-stage pseudo-labels may be less reliable; replacing them over time helps improve augmentation quality as training progresses.
3.2 Semantic-Aware Loss
To further mitigate class bias, we introduce a weighted parameter into the classification loss, guided by the Relation Contextual Module (RCM) for foreground classes. This weighting scheme emphasizes classes that are frequently misclassified as dominant majority classes, thereby directing the model’s attention toward improving performance on underrepresented or confused categories.
To accentuate this focus, we apply a non-linear transformation to the RCM values as follows:
$w_{i}=\begin{cases}\sqrt{1-\text{{RCM}}(c_{i},x_{i})},&\text{if }c_{i}=x_{i}\\ \sqrt{\frac{\text{{RCM}}(c_{i},x_{i})}{\text{{RCM}}(c_{i},c_{i})}},&\text{otherwise}\end{cases}$ (7)
where $w_{i}$ denotes the weight for the $i^{\text{th}}$ instance, $c_{i}$ is its pseudo-label, and $x_{i}$ is the predicted class. When $c_{i}\neq x_{i}$, we normalize using the diagonal entry $\text{{RCM}}(c_{i},c_{i})$ to scale the misclassification likelihood relative to the class’s overall performance.
Background class weights are uniformly set to 1 to prevent model bias against them. To reconcile the disparity between foreground and background weights, we normalize foreground instance weights so that their mean equals that of the background class:
$W_{f}=\frac{W_{f}}{\text{mean}(W_{f})},$ (8)
where $W_{f}$ represents the set of weights for foreground instances.
To avoid excessive influence from extreme weights, we apply a regularization term $\lambda_{l}$ to all class-relation weights:
$W=\frac{W+\lambda_{l}}{1+\lambda_{l}},$ (9)
ensuring smoother gradients and preventing the loss from being overly sensitive to outlier weights.
The weighted classification loss is then defined as:
$\mathcal{L}_{\text{cls}}=\frac{1}{N}\sum_{i=1}^{N}w_{i}\cdot\text{{CE}}(c_{i},x_{i}),$ (10)
where $N$ is the total number of instances, and $\text{{CE}}(\cdot)$ denotes the standard cross-entropy loss.
3.3 Grounded Supervision
We introduce a framework that leverages Large Vision Foundation Models (LVFMs) through expert distillation, as depicted in 2. We introduce an Expert branch aimed at assimilating the strengths of LVFMs into a unified learning approach. As an initial assumption, we posit that the expert model can effectively represent a wide variety of images sourced from datasets like ImageNet1K or ImageNet21k [26], LAION-400M [103], or DataComp1B [34]. We have considered many prominent families of teacher models: CLIP [90], DINOV2 [87], ViT [32]and SAM [64], GroundingDINO [78], and BioMedParse [146]. These models have exhibited outstanding performance across a wide array of tasks, with CLIP demonstrating broad proficiency, DINOV2 excelling in downstream dense tasks such as semantic segmentation under linear probe, and SAM particularly shining in open-vocabulary segmentation tasks. The newly introduced Expert branch incorporates the pretrained LVFMs. In the context of knowledge distillation, utilizing a single expert’s embedding directly for distilling knowledge to the student model may not suffice. We found that relying on an expert’s embedding directly for distilling knowledge to the student model is insufficient for effective pseudo-label correction during adaptation. Hence, we propose using expert predictions and introducing a pseudo-supervised loss to guide the student model out from the local minima.
Our objective is to transfer the expertise of the expert model to the student model. The expert model $E(\cdot)$ is frozen, while the student model $S(\cdot)$ is updated via gradient descent during adaptation. To supervise the student using pseudo labels from the expert, we employ a combination of a consistency loss (mean teacher loss) and a bounding box regression loss.
Let $x_{i}$ be an input image. The expert produces pseudo-labels in the form of bounding box predictions $E_{\text{bbox}}^{i}=E(x_{i})$, which are used to supervise the student’s prediction $S_{\text{bbox}}^{i}=S(x_{i})$. We define the loss as:
$\mathcal{L}_{\text{expert}}=$ $\lambda_{\text{cls}}\cdot\mathcal{L}_{\text{cls}}(S_{\text{bbox}}^{i},E_{\text{bbox}}^{i})$
$+\lambda_{\text{reg}}\cdot\mathcal{L}_{\text{reg}}(S_{\text{bbox}}^{i},E_{\text{bbox}}^{i})$ (11)
where $\mathcal{L}_{\text{cls}}$ is the classification loss (semantic aware loss), and $\mathcal{L}_{\text{reg}}$ is the bounding box regression loss (e.g., smooth L1 or GIoU loss). The weights $\lambda_{\text{cls}}$ and $\lambda_{\text{reg}}$ control the contribution of each term.
This unified loss ensures that the student model learns both semantic consistency and localization accuracy from the expert model via pseudo-supervision.
3.4 Overall Objective and Training Strategy.
The final objective combines supervised, unsupervised, and discriminator losses:
$\mathcal{L}=\lambda_{u}\mathcal{L}_{\text{stu}}+\lambda_{d}\mathcal{L}_{\text{dis}}+\mathcal{L}_{\text{Expert}},$ (12)
where $\lambda_{u}$ and $\lambda_{d}$ control the contributions of the unsupervised and discriminator losses, respectively, and $\mathcal{L}_{\text{stu}},\ \text{and}\ \mathcal{L}_{\text{Expert}}$ have been formulated in Eq. (1), and (11) respectively. As the teaching process continues, an adequate amount of pseudo-boxes will be produced, and the influence of $\lambda_{u}$ should be lowered to avoid the encoder over-fitting to pseudo-labels. As a result, We decay $\lambda_{d}$ as adaptation continues. More details are discussed in the Supplementary material.
4 Experiments
We conduct extensive experiments to assess the effectiveness of our approach across various challenging medical imaging datasets, as well as standard UDA and SFOD benchmarks. In the UDA setting, both source and target domain data are accessible during training, facilitating adaptation. Conversely, in the SFOD scenario, adaptation is performed using only a source-trained model without access to source domain data. Additionally, we perform ablation studies employing diverse exchange strategies to validate the efficacy of our proposed methods. We further analyze the promising results obtained by our method through comprehensive visualizations and component-wise evaluations.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
UMT [137] CVPR’21 0.01 0.09 0.15 0.21 0.204 0.319
SFA [114] MM’21 0.03 0.13 0.23 0.35 0.241 0.338
H2FA [126] CVPR’22 0.13 0.32 0.45 0.52 0.575 0.312
AQT [53] IJCAI’22 0.11 0.37 0.44 0.57 0.680 0.349
AT [71] CVPR’22 0.19 0.38 0.65 0.75 0.721 0.512
DDSM D-Adapt [55] ICLR’22 0.11 0.29 0.45 0.59 0.731 0.414
to HT [29] CVPR’23 0.17 0.49 0.61 0.69 0.704 0.362
INBreast CLIPGAP [109] CVPR’23 0.15 0.55 0.61 0.75 0.712 0.445
ConfMIX [84] WACV’23 0.19 0.47 0.58 0.73 0.737 0.409
MRT [148] ICCV’23 0.16 0.54 0.64 0.72 0.789 0.489
D-MASTER [6] MICCAI’24 0.25 0.61 0.70 0.82 0.808 0.524
Mexforer [114] MM’21 0.02 0.03 0.03 0.03 0.060 0.090
IRG [27] CVPR’23 0.05 0.05 0.07 0.09 0.110 0.120
GT (Ours) 0.06 0.45 0.65 0.92 0.589 0.758
Table 1: Results of adaptation from DDSM to INBreast.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
D-Adapt [55] ICLR’21 0.04 0.12 0.18 0.29 0.439 0.263
AT [73] CVPR’22 0.16 0.28 0.35 0.42 0.486 0.338
DDSM H2FA [127] CVPR’22 0.03 0.13 0.18 0.36 0.634 0.236
to MRT [147] ICCV’23 0.32 0.52 0.69 0.72 0.741 0.352
RSNA Mexforer [114] MM’21 0.24 0.31 0.39 0.39 0.336 0.287
IRG [27] CVPR’23 0.16 0.25 0.37 0.39 0.308 0.235
GT (Ours) 0.10 0.43 0.58 0.91 0.781 0.530
Table 2: Results of adaptation from DDSM to RSNA.
Exp Method Venue SF R@0.05 R@0.3 R@0.5 R@1.0 AUC F1-score
D-Adapt [55] ICLR’21 0.00 0.06 0.09 0.10 0.381 0.362
AT [73] CVPR’22 0.01 0.08 0.10 0.15 0.385 0.311
RSNA H2FA [127] CVPR’22 0.02 0.08 0.10 0.12 0.483 0.315
to MRT [147] ICCV’23 0.03 0.09 0.12 0.17 0.739 0.587
INBreast Mexforer [114] MM’21 0.02 0.03 0.03 0.03 0.060 0.090
IRG [27] CVPR’23 0.05 0.05 0.07 0.09 0.110 0.120
GT (Ours) 0.01 0.28 0.49 0.90 0.638 0.605
Table 3: Results of adaptation from RSNA to INBreast.
Method Venue SF Person Rider Car Truck Bus Train Mcycle Bicycle mAP
DA-Faster [18] CVPR’18 29.2 40.4 43.4 19.7 38.3 28.5 23.7 32.7 32.0
EPM [49] ECCV’20 44.2 46.6 58.5 24.8 45.2 29.1 28.6 34.6 39.0
SSAL [86] NIPS’21 45.1 47.4 59.4 24.5 50.0 25.7 26.0 38.7 39.6
SFA [114] MM’21 46.5 48.6 62.6 25.1 46.2 29.4 28.3 44.0 41.3
UMT [28] CVPR’21 33.0 46.7 48.6 34.1 56.5 46.8 30.4 37.3 41.7
D-adapt [55] ICLR’21 40.8 47.1 57.5 33.5 46.9 41.4 33.6 43.0 43.0
TIA [144] CVPR’22 34.8 46.3 49.7 31.1 52.1 48.6 37.7 38.1 42.3
PT [44] ICML’22 40.2 48.8 63.4 30.7 51.8 30.6 35.4 44.5 42.7
MTTrans [136] ECCV’22 47.7 49.9 65.2 25.8 45.9 33.8 32.6 46.5 43.4
SIGMA [68] CVPR’22 44.0 43.9 60.3 31.6 50.4 51.5 31.7 40.6 44.2
O2net [41] MM’22 48.7 51.5 63.6 31.1 47.6 47.8 38.0 45.9 46.8
AQT [53] IJCAI’22 49.3 52.3 64.4 27.7 53.7 46.5 36.0 46.4 47.1
AT [71] CVPR’22 43.7 54.1 62.3 31.9 54.4 49.3 35.2 47.9 47.4
TDD [45] CVPR’22 50.7 53.7 68.2 35.1 53.0 45.1 38.9 49.1 49.2
CIGAR [79] CVPR’23 45.3 45.3 61.6 32.1 50.0 51.0 31.9 40.4 44.7
CSDA [38] ICCV’23 46.6 46.3 63.1 28.1 56.3 53.7 33.1 39.1 45.8
HT [30] CVPR’23 52.1 55.8 67.5 32.7 55.9 49.1 40.1 50.3 50.4
MRT [147] ICCV’23 52.8 51.7 68.7 35.9 58.1 54.5 41.0 47.1 51.2
SFOD [69] AAAI’21 21.7 44.0 40.4 32.6 11.8 25.3 34.5 34.3 30.6
SFOD-Mosaic [69] AAAI’21 25.5 44.5 40.7 33.2 22.2 28.4 34.1 39.0 33.5
HCL [52] NIPS’21 38.7 46.0 47.9 33.0 45.7 38.9 32.8 34.9 39.7
SOAP [124] IJIS’21 35.9 45.0 48.4 23.9 37.2 24.3 31.8 37.9 35.5
LODS [67] CVPR’22 34.0 45.7 48.8 27.3 39.7 19.6 33.2 37.8 35.8
AASFOD [21] AAAI’23 32.3 44.1 44.6 28.1 34.3 29.0 31.8 38.9 35.4
IRG [110] CVPR’23 37.4 45.2 51.9 24.4 39.6 25.2 31.5 41.6 37.1
PETS [77] ICCV’23 46.1 52.8 63.4 21.8 46.7 5.5 37.4 48.4 40.3
DACA [145] IJCV’24 44.7 31.2 60.1 53.1 53.9 0.0 27.8 45.9 39.9
LPLD [133] ECCV’24 39.7 49.1 56.6 29.6 46.3 26.4 36.1 43.6 40.9
SF-UT [43] ECCV’24 40.9 48.0 58.9 29.6 51.9 50.2 36.2 44.1 45.0
GT (Ours) 42.7 55.4 61.7 40.7 62.0 54.6 39.1 53.0 50.8
Oracle 66.3 61.1 80.8 45.6 68.8 52.0 49.1 54.9 59.8
Table 4: Results of adaptation from Normal Cityscapes to Foggy weather Cityscapes (CF). “SF” refers to source-free setting. “Oracle” refers to the models trained by using labels during training.
4.1 Comparison with Current SOTA Methods
We evaluate the performance of our proposed Grounded Teacher (GT) method against other approaches on all three medical benchmarks and the natural benchmark mentioned earlier for generalizability. Since UDA and SFDOAD share similar task settings, we conducted comparisons with both. Table 1, Table 2 and Table 3 present the comparison results on medical image datasets. Table 4 presents comparison of the natural dataset. Our proposed GT consistently outperforms existing all SOTA methods, demonstrating generalizability and significant improvements across both natural and medical settings.
DDSM to INBreast. Adaptation from large to small-scale medical datasets with different modalities. Here, we consider DDSM [66] dataset as the source domain and INBreast [85] as the target domain. Results are presented in Table 1. Our proposed method demonstrates superior performance across various False positives per Image (FPI) values compared to existing methods as displayed.
RSNA to INBreast. Adaptation across medical datasets with different machine-acquisitions. This is vital for enhancing healthcare outcomes, improving diagnostic accuracy, and facilitating better clinical decisions. To evaluate our method’s performance, we adapt a model trained on DDSM [66] to RSNA [11]. Results for all FPI values are presented in Table 2, demonstrating that our method achieves state-of-the-art performance on this benchmark.
DDSM to RSNA. This experiment evaluates domain adaptation across datasets collected from distinct geographical regions. Specifically, we consider the DDSM [66] dataset as the source domain and RSNA [11] as the target domain. As shown in Table 2, our proposed method consistently outperforms existing approaches across all FPI thresholds, demonstrating its robustness and effectiveness in cross-domain generalization.
Cityscapes to Foggy Cityscapes. Object detectors often experience a significant drop in performance when deployed under adverse real-world conditions such as fog, due to the domain shift caused by the lack of such conditions in the training data. Domain adaptation aims to bridge this gap between normal and adverse weather scenarios. To investigate this, we conduct experiments on the widely-used Cityscapes $\rightarrow$ FoggyCityscapes benchmark. As presented in Table 4, student-teacher based frameworks consistently outperform non-student-teacher approaches by a notable margin. Our method, GT, built upon mean teacher architecture [71], achieves a substantial improvement with a performance of 50.8 mAP. Furthermore, it demonstrates notable gains in minority object classes, highlighting its robustness and adaptability in challenging weather conditions.
4.2 Ablation Studies
Method RCM SA SAL mAP
Base (AT [72]) 30.4
+ SA 41.8
+ SAL 44.2
GT (Full) 50.8
Table 5: Ablation studies on GT components. RCM is included in all studies as it forms the basis of CRA and SAL. We report the mean average precision at 0.50 IoU (mAP). Our contributions are not included in the base framework (AT [71]).
To verify the significance of our contributions, we conducted an ablation study. All experiments within this study were performed on the Cityscapes $\to$ Foggy Cityscapes benchmark using the VGG16 backbone.
4.2.1 Quantitative Ablation
Table 5 presents a quantitative analysis of the impact of each component within our framework. The base framework prior to integrating our modules aligns with the AT model [71], achieving a 30.4 mAP. Given that the Relation-Contextual Module (RCM) is integral to both class-relation augmentation and semantic-aware loss, it remains a constant across all experimental variations. Integrating Semantic Augmentation (SA) along with RCM increases the performance to 41.8 mAP, a gain of +11.4 over the base. Adding the Semantic Aware Loss (SAL) with RCM separately results in a higher mAP of 44.2, demonstrating a +13.8 improvement compared to the baseline, indicating that SAL provides a greater performance boost than SA in this ablation setting. Notably, Class-Relation Augmentation (CRA) significantly reduces the performance disparity between minority and majority classes, as evidenced. Finally, our full GT model, which combines RCM, SA, and SAL, achieves the highest score of 50.8 mAP, which not only boosts overall performance but also underscores our method’s efficacy in managing class imbalance.
Figure 3: FROC analysis of model generalization on unseen target domains. Performance comparison between our method (GT) and the MT baseline for detection tasks under domain shift: (a) DDSM→INBreast (410 images), (b) DDSM→RSNA (993 images), and (c) RSNA→INBreast (410 images), where our method demonstrates superior performance across different domain shifts. Higher curves indicate better trade-offs between sensitivity (TPR) and false positives per image (FPI).
Figure 4: Qualitative results of our proposed Grounded Teacher on Medical settings (DDSM $\rightarrow$ INBreast). We present visual comparisons between AT (top row) and GT (bottom row). Our method, GT, effectively identifies Malignancies better and mitigates false negatives. Bounding box colors indicate: Blue — Ground truth, Green — true positive, and Red — false negative.
Figure 5: Qualitative results of our proposed Grounded Teacher on Natural settings (Cityscapes $\rightarrow$ Foggy Cityscapes). We present visual comparisons between AT (top row) and GT (bottom row). Our method, GT, effectively mitigates misclassifications, reduces false negatives, and eliminates false positives. Bounding box colors indicate: Green — true positive, Blue — misclassified, Red — false negative, and Pink — false positive.
4.2.2 Impact of Augmentation
In our study, we assess the influence of augmentation strategies, focusing on both the proportion of augmented data and the criteria for selecting instances to augment. Image augmentation plays a pivotal role in our methodology, enhancing the dataset by incorporating additional representations, particularly of minority classes. However, excessive augmentation can lead to the model learning less accurate class representations. To mitigate this, we judiciously augment a random subset of images.
The pairing of class instances during augmentation is crucial for improving its effectiveness. Unlike random pairing, our Semantic Augmentation (SA) method prioritizes pairing instances from closely related minority and majority classes. This targeted approach ensures that the Mixup outputs are more meaningful, thereby enhancing the performance of minority classes.
Table 6 presents the experimental results of this approach. We compare the effects of randomly selecting class instances for Mixup against our SA method across different augmentation probabilities, which represent the likelihood of an instance in the base image being augmented. Applying Mixup randomly yields a modest improvement of +1.1 mAP in overall performance. In contrast, employing SA leads to a more substantial increase of +4.4 mAP. These findings indicate that pairing highly related classes during Mixup enhances minority class performance with minimal impact on majority classes. Furthermore, our experiments reveal that excessive augmentation can negatively affect model performance, underscoring the importance of a balanced augmentation strategy.
Selection Method mAP
Random (0.5) 46.4
Semantic Augmentation (1.0) 47.5
Semantic Augmentation (0.5) 50.8
Table 6: Comparison of selecting class instances randomly and via CRA. Values in brackets refer to the likelihood of an instance being augmented.
4.2.3 Weighing Strategy for Semantic Loss
To enhance the performance of minority classes, we introduce a Semantic Aware Loss (SAL) function, as defined in Eq 10. Class-level loss functions are a prevalent strategy for addressing dataset class imbalances. We compare our SAL with a variant of existing class-level losses that utilize only the diagonal elements of the Inter-Class Relation module, which correspond to the ground-truth class likelihood for accurate classification. In contrast, SAL leverages the likelihood of the ground-truth being classified as the predicted class to influence its loss function.
As shown in Table 7, incorporating inter-class relationships through SAL achieves an mAP of 50.8, a performance improvement of +3.7 mAP compared to the base. While SAL effectively utilizes inter-class information, there are instances where it might excessively penalize well-performing classes without proper constraints. To mitigate this, we apply a regularization term, as specified in Eq 9. Omitting this regularization (SAL w/o Reg.) results in an mAP of 47.8, demonstrating a significant performance drop of -3.0 mAP compared to the full SAL. This clearly highlights the importance of the regularization term, likely preventing certain class weights from becoming disproportionately small during training.
Class Weight Strategy mAP
Class-Level 47.1
SAL w/o Reg. 47.8
SAL 50.8
Table 7: Class loss weighting strategies. Class-Level uses only the diagonal values in our RCM with regularization. SAL refers to our full Semantic Aware Loss.
4.2.4 Qualitative Results
Figure 5 presents the qualitative outcomes of our method, comparing predictions from baseline AT [71] in the top row with those from the Ground Truth (GT) in the bottom row. The GT effectively corrects misclassifications, as highlighted by the blue boxes. For instance, in columns 1, 2, and 4, the GT accurately labels ’person’, ’car’, and ’truck’, respectively, rectifying the AT’s errors. Moreover, the GT demonstrates a notable reduction in both false positives and false negatives, indicated by the pink and red boxes, respectively. This enhancement signifies improved detection accuracy across various object scales and classes. This improvement is particularly evident in column 3, where the GT successfully identifies small-scale objects and provides more accurate detections for larger-scale objects. Also in Fig 5, we show how qualitatively the predictions are improved with various proposed modules.